# USD: Unsupervised Soft Contrastive Learning for Fault Detection in Multivariate Time Series

### v1

@@ V Experiments
Datasets information. We selected the five most commonly used datasets for fault detection to evaluate the effectiveness of our method. These datasets are widely recognized in the field of One-Class Classification (OCC) for Multivariate Time Series (MTS) anomaly detection. The datasets include:
@@
  • SMD (Server Machine Dataset) [24]: Collected from a large internet company’s server farms, this dataset comprises metrics from different server machines, providing a basis for detecting unusual behaviors in server operations  [^5] https://github.com/NetManAIOps/OmniAnomaly.
TABLE I: Dataset information. The number of metrics, training size, test size, training anomaly ratio, and test anomaly ratio of each dataset.
Dataset
Number
of
Metric
Training
Size
Test
Size
Training
Anomaly
Ratio (%)
Training
Anomaly
Ratio (%)
SWaT 51 269,951 89,984 17.7 5.2
WADI 123 103,680 69,121 6.4 4.6
PSM 25 52,704 35,137 23.1 34.6
MSL 55 44,237 29,492 14.7 4.3
SMD 38 425,052 283,368 4.2 4.1
TABLE II: Fault detection performance of Area Under the Receiver Operating Characteristic (AUROC) on five public datasets. We compare the performance of the MTGFLOW method using a consistent window size (USD) and our optimal results with varying window sizes (USD*).The best results are highlighted in bold.
SWaT WADI PSM MSL SMD
DROCC [10] ICML, 2020 72.6(±3.8) 75.6(±1.6) 74.3(±2.0) 53.4(±1.6) 76.7(±8.7) 70.5 6
DeepSAD [23] ICLR, 2020 75.4(±2.4) 85.4(±2.7) 73.2(±3.3) 61.6(±0.6) 85.9(±11.1) 76.3 5
USAD [4] KDD, 2020 78.8(±1.0) 86.1(±0.9) 78.0(±0.2) 57.0(±0.1) 86.9(±11.7) 77.4 4
DAGMM [36] ICLR, 2018 72.8(±3.0) 77.2(±0.9) 64.6(±2.6) 56.5(±2.6) 78.0(±9.2) 69.8 7
GANF [8] ICLR, 2022 79.8(±0.7) 90.3(±1.0) 81.8(±1.5) 64.5(±1.9) 89.2(±7.8) 81.1 3
MTGFlow [34] AAAI, 2023 84.8(±1.5) 91.9(±1.1) 85.7(±1.5) 67.2(±1.7) 91.3(±7.6) 84.2 2
MTGFlow_Cluster [35] TKDE, 2024 83.1(±1.3) 91.8(±0.4) 87.1(±2.4) 68.2(±2.6) - - -
USD OURS 90.2(±0.9) 94.1(±0.4) 88.9(±2.4) 74.2(±2.0) 94.2(±6.5) 88.3 1
USD* OURS 92.0(±1.2) 94.9(±1.2) 88.9(±2.4) 76.0(±1.1) 94.2(±6.5) 89.2 1*
TABLE III: Anomaly detection performance of Precision-Recall(PR) on five public datasets. We compare the performance of the MTGFLOW method using a consistent window size (USD) and our optimal results with varying window sizes (USD*).The best results are highlighted in bold.
SWaT WADI PSM MSL SMD
DROCC [10] ICML, 2020 26.4(±9.8) 16.7(±7.1) 60.7(±11.4) 13.2(±0.9) 20.7(±15.6) 27.5 6
DeepSAD [23] ICLR, 2020 45.7(±12.3) 23.1(±6.3) 66.7(±10.8) 26.3(±1.7) 61.8(±21.4) 44.7 4
USAD [4] KDD, 2020 18.8(±0.6) 19.8(±0.5) 57.9(±3.6) 31.3(±0.0) 68.4(±19.9) 39.2 5
GANF [8] ICLR, 2022 21.6(±1.8) 39.0(±3.1) 73.8(±4.7) 31.1(±0.2) 64.4(±21.4) 46.0 3
MTGFlow [34] AAAI, 2023 38.6(±6.1) 42.2(±4.9) 76.2(±4.8) 31.1(±2.6) 64.2(±21.5) 50.5 2
USD OURS 53.4(±10.5) 45.5(±4.9) 83.2(±3.9) 31.9(±3.8) 73.6(±19.7) 57.5 1
USD* OURS 59.1(±5.5) 60.3(±5.5) 83.2(±3.9) 45.9(±1.5) 73.6(±19.7) 64.4 1*
Implementation details. For all datasets, we set the window size to 60 and the stride size to 10. Additional specific parameters can be found in Appendix B. Model Hyperparameter Settings and Model Hyperparameter Settings. All experiments were run for 400 epochs and executed using PyTorch 2.2.1 on an NVIDIA RTX 3090 24GB GPU.
V-B Evaluation Metric and Baselines Methods.
Consistent with prior research, our Unsupervised Anomaly Detection (USD) method is designed to identify window-level anomalies. Anomalies are labeled at the window level if any point within the window is anomalous. To evaluate the performance of our approach, we utilize two key metrics:
Area Under the Receiver Operating Characteristic Curve (AUROC): This metric helps in assessing the overall effectiveness of the model in distinguishing between the normal and anomalous windows across various threshold settings.
Area Under the Precision-Recall Curve (AUPRC): Given the potential class imbalance in anomaly detection scenarios, the AUPRC is particularly useful. It provides a measure of the model’s ability to capture true anomalies (precision) while minimizing false positives, across different decision thresholds.
These metrics allow us to comprehensively evaluate the detection capabilities of our model under different operational conditions and threshold levels.
Baselines. We compare our method, MTGFlow, against state-of-the-art (SOTA) semi-supervised and unsupervised methods. The semi-supervised methods include:
@@
The unsupervised methods include:
DROCC [10]: A robust one-class classification method applicable to most standard domains without requiring additional information. It assumes that points from the class of interest lie on a well-sampled, locally linear low-dimensional manifold, effectively avoiding the issue of representation collapse
@@
DAGMM [36]: An unsupervised anomaly detection method that generates low-dimensional representations and reconstruction errors using a deep autoencoder, which are then input into a Gaussian Mixture Model (GMM). DAGMM optimizes the parameters of both the autoencoder and the mixture model in an end-to-end manner, achieving efficient anomaly detection.
@@
We evaluate the performance degradation of semi-supervised methods when dealing with contaminated training sets and compare our USD with SOTA unsupervised density estimation methods.
V-C Fault Detection Performance on Five Benchmark Dataset
First, we discuss the performance advantages of the USD method by comparing results on five common benchmarks. We list the results for the AUROC and AUPRC metrics in Tables II and  III, where the numbers in parentheses represent the variance from five different seed experiments. We report on two variants of USD: one (USD) uses the same hyperparameters across all experiments, and the other (USD*) involves hyperparameter tuning specific to each target dataset. The best results are highlighted in bold. To visualize the benefits of USD performance, the ROC and PR curves on four dataset is showned in Fig. 3. In the ROC and PR curves, the USD method consistently outperforms the baseline methods, demonstrating its robustness and effectiveness in anomaly detection tasks.
Performance improvement. In terms of performance improvements, as shown in Tables II, the USD method shows significant enhancements compared to state-of-the-art methods. These improvements are evident across both PR and AUC curves, particularly on the SWAT dataset with a 5.4% increase in AUROC and a 7.7% increase in AUPRC, and on the MSL dataset with a 6% increase in AUROC and a 7.0% increase in AUPRC. On average, there is a 4.1% increase in AUROC performance and a 7.0% increase in AUPRC performance. The consistent improvements in both AUROC and AUPRC indicate that the USD method has enhanced performance across multiple aspects.
Stability improvement. As shown in Tables II and  III and Fig. 3, the stability of the USD method is also noteworthy, exhibiting lower variance across all test datasets compared to traditional methods. This indicates that the USD method is more robust and less sensitive to random initialization, making it a reliable choice for anomaly detection tasks.
Performance potential. Furthermore, by comparing the results of USD and USD*, we find that conducting hyperparameter searches for each dataset can significantly boost model performance. This suggests that our method has good hyperparameter adaptability and considerable potential for further performance enhancements.
Figure 4: Violin plot of AUROC scores on different training splits. The x-axis represents the training split ratio, and the y-axis represents the AUROC scores. The USD method consistently outperforms the baseline methods(MTGFLOW) across different training splits, demonstrating its robustness and effectiveness in anomaly detection tasks.
Additional results on difference train/validation splits. To further investigate the influence of anomaly contamination rates, we vary training splits to adjust anomalous contamination rates. For all the above-mentioned datasets, the training split increases from 60% to 80% with 5% stride. We present an average result over five runs in Fig. 4. Although the anomaly contamination ratio of training dataset rises, the anomaly detection performance of MTGFlow remains at a stable high level. This indicates that the proposed USD method is robust to the anomaly contamination ratio of the training dataset.
V-D Visualization of multi-subclass data distributions
Figure 5: Visualization of multi-subclass data distributions.(a) The t-SNE visualization of embedding of USD on SWAT datasets. Multiple clusters can be seen in the representation, which is consistent with the multiple manifold assumption of this paper. The normal and abnormal states are not a single cluster, but can be divided into separate subclusters. (b) The heatmaps of individual data point, each row in the heatmap is the index number of the time and each column is the index number of the sensor. We selected four sub-clusters, and two samples from each cluster were randomly selected for presentation. The similarity of samples within sub-clusters and the difference of samples between sub-clusters are illustrated.
Figure 6: Point plot of the anomalies predict outputs of MTGFLOW and proposed USD. The x-axis represents the anomaly index, and the y-axis represents the log-likelihood of the anomaly. Anomalous ground truths are marked by a red background. This indicates that USD is able to detect anomalies more accurately and in a timely manner.
To further investigate the effectiveness of the USD method, we visualize the multi-subclass data distributions on the SWAT dataset. As shown in Fig. 5, the t-SNE visualization of the embedding of USD on SWAT datasets reveals multiple clusters, consistent with the multiple manifold assumption of this paper. The normal and abnormal states are not a single cluster but can be divided into separate subclusters. The heatmaps of individual data points further illustrate the similarity of samples within sub-clusters and the differences between samples in different sub-clusters. This visualization demonstrates the ability of the USD method to capture the diverse patterns present in both normal and abnormal states, enhancing its anomaly detection capabilities.
In addition, we present the point plot of the anomalies predict outputs of MTGFLOW and the proposed USD in Fig. 6. The x-axis represents the anomaly index, and the y-axis represents the log-likelihood of the anomaly. Anomalous ground truths are marked by a red background. This indicates that USD is able to detect anomalies more accurately and in a timely manner, outperforming MTGFLOW in terms of anomaly detection performance.
Fig. 7 shows the bar plot of the frequency of normalized anomaly scores on baseline methods and USD. The x-axis represents the anomaly scores, and the y-axis represents the frequency of each score. The normalized anomaly scores of USD are significantly lower than those of GANF and MTGFlow, indicating that USD is more effective at distinguishing between normal and abnormal states.
Figure 7: The bar plot of the frequency of normalized anomaly scores on baseline methods and USD. The x-axis represents the anomaly scores, and the y-axis represents the frequency of each score. The normalized anomaly scores of USD are significantly lower than those of GANF and MTGFlow, indicating that USD is more effective at distinguishing between normal and abnormal states.
V-E Ablation Study: The Effect of SCL Loss.
TABLE IV: Ablation study. AUROC comparison on four datasets (SWaT, WADI, PSM , and MSL). The best results are in bold. The results are averaged over 5 runs, and the standard deviation is shown in parentheses.
Method SWaT WADI PSM MSL
USD 90.2(±0.9) 94.1(±0.4) 88.9(±2.4) 74.2(±2.0)
w. Gaussian 88.9(±1.1) 93.9(±1.7) 86.7(±2.5) 72.1(±0.7)
w/o. SCL 84.3(±1.5) 91.7(±1.2) 85.5(±1.3) 70.7(±2.4)
w/o. CL 84.7(±1.3) 90.8(±1.4) 85.9(±1.6) 69.2(±2.1)
w/o. Aug
84.8(±1.5) 91.9(±1.1) 85.7(±1.5) 67.2(±1.7)
To assess the effectiveness of each component designed in our model, we conducted a series of ablation experiments. The results of these experiments are presented in Table IV.
We performed controlled experiments to verify the necessity of the Soft Contrastive Learning (SCL) Loss in the UNDA settings across all four datasets. The performance of the proposed USD method is denoted as USD in the table. The variant w/o. SCL represents the model performance with the SCL loss component, $\mathcal{L}_{SCL}(\mathbf{x}_{i}^{t},{y_{i}}^{t})$, removed from the overall loss function of USD. The variant w. CL indicates the model performance when the SCL loss is replaced by a typical Contrastive Loss (CL), $\mathcal{L}_{\text{CL}}$, as defined in Eq. (10).
The results clearly demonstrate that the SCL Loss significantly outperforms the traditional CL loss. We attribute the inferior performance of $\mathcal{L}_{\text{CL}}$ to its inability to adequately address the view-noise caused by domain bias. Conversely, the SCL loss is more effective at mitigating this issue, leading to better model performance across the datasets.
V-F Hyperparameter Robustness

### v2

@@ A. Details of Dataset
We selected the five most commonly used datasets for fault detection to evaluate the effectiveness of our method (in Table VIII). These datasets are widely recognized in the field of One-Class Classification (OCC) for Multivariate Time Series (MTS) anomaly detection. The datasets include:
@@
  • SMD (Server Machine Dataset) [27]: Collected from a large internet company’s server farms, this dataset comprises metrics from different server machines, providing a basis for detecting unusual behaviors in server operations  [^5] https://github.com/NetManAIOps/OmniAnomaly. For this dataset, the average results across all entities are reported. For consistency and proper evaluation, we adjusted the random seed during dataset splitting to ensure anomalies are present in the test set while maintaining a training/testing ratio.
TABLE VIII: Dataset information. The number of metrics, training size, test size, training anomaly ratio, and test anomaly ratio of each dataset.
Property SWaT WADI PSM MSL SMD
# of Metrics 51 123 25 55 38
Training Size 269,951 103,680 52,704 44,237 425,052
Test Size 89,984 69,121 35,137 29,492 283,368
Training & Test Size 359,935 172,801 87,841 73,729 708,420
Training Anomaly Ratio (%) 17.7 6.4 23.1 14.7 4.2
Test Anomaly Ratio (%) 5.2 4.6 34.6 4.3 4.1
A. Details of Baselines
Baselines. We compare our method, DMTFD, against state-of-the-art (SOTA) methods.
These methods include: DROCC [12]: A robust one-class classification method applicable to most standard domains without requiring additional information. It assumes that points from the class of interest lie on a well-sampled, locally linear low-dimensional manifold, effectively avoiding the issue of representation collapse
@@ A. Details of Baselines
We evaluate the performance degradation of semi-supervised methods when dealing with contaminated training sets and compare our DMTFD with SOTA unsupervised density estimation methods.
A. Details of SoftCL loss
A.1 Details of the transformation from Eq. (6) to Eq. (7)
A.2 The proposed SoftCL loss is a smoother CL loss
This proof tries to indicate that the proposed SoftCL loss is a smoother CL loss. We discuss the differences by comparing the two losses to prove this point. the forward propagation of the network is, ${z}_{i}=H(\hat{z}_{i}),\hat{z}_{i}=F(x_{i})$, ${z}_{j}=H(\hat{z}_{j}),\hat{z}_{j}=F(x_{j})$. We found that we mix $y$ and $\hat{z}$ in the main text, and we will correct this in the new version. So, in this section ${z}_{i}=H(y_{i}),y_{i}=F(x_{i})$, ${z}_{j}=H(y_{j}),y_{j}=F(x_{j})$ is also correct.
Let $H(\cdot)$ satisfy $K$-Lipschitz continuity, then $d^{z}_{ij}=k^{*}d^{y}_{ij},k^{*}\in[1/K,K],$ where $k^{*}$ is a Lipschitz constant. The difference between $L_{\text{SoftCL}}$ loss and $L_{\text{CL}}$ loss is,
$L_{\text{CL}}-L_{\text{SoftCL}}\approx\sum_{j}\biggl{[}\left(\mathcal{H}_{ij}-[1+(e^{\alpha}-1)\mathcal{H}_{ij}]\kappa\left(d^{y}_{ij}\right)\right)\log\left(\frac{1}{\kappa\left(d_{ij}^{z}\right)}-1\right)\biggl{]}.$ (12)
Because the $\alpha>0$, the proposed SoftCL loss is the soft version of the CL loss. if $\mathcal{H}_{ij}=1$, we have:
$(L_{\text{CL}}-L_{\text{SoftCL}})|_{\mathcal{H}_{ij}=1}=\sum\biggl{[}\left((1-e^{\alpha})\kappa\left(k^{*}d^{z}_{ij}\right)\right)\log\left(\frac{1}{\kappa\left(d_{ij}^{z}\right)}-1\right)\biggl{]}$ (13)
then:
$\lim_{\alpha\to 0}(L_{\text{CL}}-L_{\text{SoftCL}})|_{\mathcal{H}_{ij}=1}=\lim_{\alpha\to 0}\sum\biggl{[}\left((1-e^{\alpha})\kappa\left(k^{*}d^{z}_{ij}\right)\right)\log\left(\frac{1}{\kappa\left(d_{ij}^{z}\right)}-1\right)\biggl{]}=0$ (14)
Based on Eq.(14), we find that if $i,j$ is neighbor ($\mathcal{H}_{ij}=1$) and $\alpha\to 0$, there is no difference between the CL loss $L_{\text{CL}}$ and SoftCL loss $L_{\text{SoftCL}}$. When if $\mathcal{H}_{ij}=0$, the difference between the loss functions will be the function of $d_{ij}^{z}$. The CL loss $L_{\text{CL}}$ only minimizes the distance between adjacent nodes and does not maintain any structural information. The proposed SoftCL loss considers the knowledge both comes from the output of the current bottleneck and data augmentation, thus less affected by view noise.
The SoftCL loss is written as:
$L_{\text{SoftCL}}$ =$=-\sum\left\{P_{ij}\log Q_{ij}+\left(1-P_{ij}\right)\log\left(1-Q_{ij}\right)\right\}$ (16)
According to Eq. (4) and Eq. (5), we have
$Q_{ij}$ $=\kappa(d_{ij}^{z})=\kappa(z_{i},z_{j}),$
The difference between the two loss functions is:
$L_{\text{CL}}-L_{\text{SoftCL}}$ (18)
$L_{\text{CL}}-L_{\text{SoftCL}}=\sum\biggl{[}\left(\mathcal{H}_{ij}-[1+(e^{\alpha}-1)\mathcal{H}_{ij}]\kappa\left(d^{y}_{ij}\right)\right)\log\left(\frac{1}{\kappa\left(d_{ij}^{z}\right)}-1\right)\biggl{]}$ (19)
We construct the inverse mapping of $H(\cdot)$ to $H^{-1}(\cdot)$,
and then there exists $k^{*}$:
Substituting the Eq.(22) into Eq.(19).
$L_{\text{CL}}-L_{\text{SoftCL}}=\sum\biggl{[}\left(\mathcal{H}_{ij}-[1+(e^{\alpha}-1)\mathcal{H}_{ij}]\kappa\left(k^{*}d^{z}_{ij}\right)\right)\log\left(\frac{1}{\kappa\left(d_{ij}^{z}\right)}-1\right)\biggl{]}$ (23)
A.3 SoftCL is better than CL in view-noise
To demonstrate that compared to contrastive learning, the proposed SoftCL Loss has better results, we first define the signal-to-noise ratio (SNR) as an evaluation metric.
$SNR=\frac{PL}{NL}$ (24)
$\frac{PL_{cl}}{NL_{cl}}<\frac{PL_{SoftCL}}{NL_{SoftCL}}$ (25)
First, we organize the ${PL_{SoftCL}}$ and ${PL_{cl}}$ into 2 cases, $H~{}(x_{i},x_{j})=1$ and $H~{}(x_{i},x_{j})=0$, for writing convenience, we write $S~{}(z_{i},z_{j})$ as $S$ and $S^{\prime}$, respectively.
$PL_{SoftCL}=-M\left\{\left(1-S^{\prime}\right)\log\left(1-S^{\prime}\right)+S^{\prime}\log S^{\prime}\right\}-\left\{\left(1-e^{\alpha}S\right)\log(1-S)+e^{\alpha}S\log S\right\}$ (26)
$PL_{cl}=-M\log\left(1-S^{\prime}\right)-\log S$ (27)
$\ PL_{SoftCL}-PL_{cl}$ (28)
Because we have bounded that $e^{\alpha}S<=1$, so we could easily get:
$(e^{\alpha}S-1)log\frac{S}{\left(1-S\right)}<=0$ (29)
Also, we could get:
So we get:
$PL_{SoftCL}-PL_{cl}>0$ (31)
And for the case of noise pair, the values of $S$ and $S^{\prime}$ are of opposite magnitude, so obviously, there is $NL_{SoftCL}-NL_{cl}<0$.
So the formula Eq. (25) has been proved.
A. Details of Proof
-A Theoretical Analysis: Generalization Bound of MML
To better understand the theoretical robustness of the proposed Multi-Manifold Loss (MML), we derive a generalization error bound based on Rademacher complexity. Our goal is to show that under the multi-manifold hypothesis and soft similarity kernel, the model achieves improved generalization.
Notation.
Let $\mathcal{X}\subset\mathbb{R}^{d}$ denote the input space, and $\mathcal{Z}$ the representation space learned by the encoder $f:\mathcal{X}\rightarrow\mathcal{Z}$. The training set $\mathcal{S}=\{(x_{i},x_{j})\}_{i,j=1}^{n}$ consists of augmented view pairs, sampled from the true data distribution $\mathcal{D}$. Let $\mathcal{L}_{\text{MML}}$ denote the Multi-Manifold Loss defined over the representation similarity between pairs.
Definition 4 (Multi-Manifold Similarity Function).
Let $\kappa^{\beta}:\mathcal{Z}\times\mathcal{Z}\to[0,1]$ be the generalized Gaussian kernel:
$\kappa^{\beta}(z_{i},z_{j})=\exp\left(-\left(\frac{\|z_{i}-z_{j}\|_{2}}{\sigma}\right)^{\beta}\right),$ (32)
where $\sigma>0$ and $\beta\in(0,2]$. This kernel controls the tail sensitivity of the similarity under the multi-manifold setting.
Definition 5 (Multi-Manifold Loss).
Given a batch of representations $\{z_{i}\}_{i=1}^{n}$, the MML loss is defined as:
$\mathcal{L}_{\text{MML}}=\frac{1}{n^{2}}\sum_{i,j}P_{i,j}\log\frac{1}{Q_{i,j}}+(1-P_{i,j})\log\frac{1}{1-Q_{i,j}},$ (33)
where $P_{i,j}$ is the softened label (similarity in input space), and $Q_{i,j}=\kappa^{\beta}(z_{i},z_{j})$ is the output similarity in representation space.
Lemma 1 (Lipschitz Continuity of $\kappa^{\beta}$).
The generalized Gaussian kernel $\kappa^{\beta}$ is Lipschitz continuous with constant $L=\frac{\beta}{\sigma^{\beta}}\exp(-1)$ for $\|z_{i}-z_{j}\|_{2}\leq\sigma$.
Proof.
We compute the derivative:
$\left|\frac{\partial\kappa^{\beta}(z_{i},z_{j})}{\partial\|z_{i}-z_{j}\|_{2}}\right|=\beta\left(\frac{\|z_{i}-z_{j}\|_{2}}{\sigma}\right)^{\beta-1}\frac{1}{\sigma}\exp\left(-\left(\frac{\|z_{i}-z_{j}\|_{2}}{\sigma}\right)^{\beta}\right).$
The maximum of this derivative is achieved when $\|z_{i}-z_{j}\|_{2}=\sigma$, yielding the upper bound $L=\frac{\beta}{\sigma^{\beta}}\exp(-1)$. ∎
Theorem 1 (Generalization Bound of MML).
Let $\mathcal{F}$ be a class of encoders $f$ such that $\|f\|_{\infty}\leq B$, and $\mathcal{L}_{\text{MML}}\circ f$ denotes the induced loss class. Then, with probability at least $1-\delta$, for any $f\in\mathcal{F}$, we have:
$\mathbb{E}_{(x_{i},x_{j})\sim\mathcal{D}}[\mathcal{L}_{\text{MML}}(f)]\leq\frac{1}{n^{2}}\sum_{i,j}\mathcal{L}_{\text{MML}}(f(x_{i},x_{j}))+2L\cdot\mathfrak{R}_{n}(\mathcal{F})+\sqrt{\frac{\log(1/\delta)}{2n}},$ (34)
where $\mathfrak{R}_{n}(\mathcal{F})$ is the Rademacher complexity of $\mathcal{F}$ and $L$ is the Lipschitz constant of the kernel $\kappa^{\beta}$.
Proof.
Follows standard Rademacher complexity-based generalization bound derivation using Lipschitz loss (see Bartlett et al., 2002). Given the boundedness of $\kappa^{\beta}$, and its Lipschitz continuity, the loss function inherits Lipschitz smoothness, enabling standard concentration bounds. ∎
Compared to traditional contrastive losses based on hard binary labels and cosine kernels, MML leverages soft labels and a tunable decay kernel, resulting in smoother gradients and improved tolerance to view-noise. The generalization bound explicitly benefits from the reduced Lipschitz constant of $\kappa^{\beta}$ for $\beta<2$ under controlled radius.
B. Model Hyperparameter Settings and Model Hyperparameter Settings
  1. 1.
    Tables IX: Hyperparameter settings for the model used in Tables I and  II.
  2. 2.
    Tables X: Hyperparameter settings for the model used in Fig.  4.
  3. 3.
    Tables XI: Hyperparameter settings for the model used in Table. XII.
TABLE IX: Describes the model parameter details in Tables I and  II when the model DMTFD calculate ROC or PR.The parameter “alpha” corresponds to $\alpha$ in Eq. (5). “batch size” refers to the number of data samples processed simultaneously during each training process. “k” means taking k neighbor points that conform to Eq. (4). The “loss weight manifold ne” and “loss weight manifold po” are the proportion of topological loss, The “lr” is the learning rate. The “n blocks” is the number of normalized flow blocks, The “name” corresponds to different datasets. The ”seed” is the random seed. The ”window size” epresents the size of the sliding window. The “train split” represents the training set division ratio.The () after the parameters indicates that the data is specific to “DMTFD*” and is different from the standard “DMTFD” parameters.The “0.6” of the train split means that 60% of the dataset is used for training and the remaining 40% is used for testing. The “0.6*”of the train split means that 60% of the dataset is used for training, 20% is used for validation, and the remaining 20% is used for testing.
Hyperparameters Datasets SWaT Wadi PSM MSL SMD
alpha 0.1 0.1 0.1 0.1 0.1
batch size 128 64 128 128 256
k 10 20 20 20 30
loss weight manifold ne 5 5 5 5 5
loss weight manifold po 1 1 1 1 5
lr 0.01 0.002 0.002 0.002 0.1
n blocks 1 5(4) 1 2 1
name SWaT Wadi PSM MSL 28 sub-datasets
window size 60(40) 60 (160) 60 60(160) 60
train split 0.6* 0.6 0.6 0.6 0.6
TABLE X: Details the specific parameters of model DMTFD in different train split ratios across various datasets, as shown in the violin plot comparison of model DMTFD in Fig.  4. The parameters used in this table have been explained in Table. IX
Hyperparameters Datasets SWaT Wadi PSM MSL SMD
alpha 0.1 0.1 0.1 0.1 0.1
batch size 128 64 128 128 256
k 10 20 20 20 30
loss weight manifold ne 5 5 5 5 5
loss weight manifold po 1 1 1 1 5
lr 0.01 0.002 0.002 0.002 0.1
n blocks 1 5 1 2 1
name SWaT Wadi PSM MSL 28 sub-datasets
train split 0.6, 0.65, 0.7, 0.75, 0.8
window size 60 60 60 60 60
TABLE XI: Describes the model parameter details in Table. XII when performing topology experiments using various sizes of sliding windows and the number of normalized flow blocks. The parameters used in this table have been explained in Table. IX
Hyperparameters Datasets SWaT Wadi PSM MSL
alpha 0.1 0.1 0.1 0.1
batch size 128 64 128 128
k 10 20 20 20
loss weight manifold ne 5 5 5 5
loss weight manifold po 1 1 1 1
lr 0.01 0.002 0.002 0.002
name SWaT Wadi PSM MSL
train plit 0.6* 0.6 0.6 0.6